# Central Dogma Transformer II An AI Microscope for Understanding Cellular Regulatory Mechanisms

### v1

@@ Discussion
The ablation study establishes a practically important principle: gene set quality, determined by cross-dataset reproducibility filtering, governs the resolution of the AI microscope. Training with 9,335 genes produced inferior predictions ($r=0.36$), while 2,361 cross-dataset validated genes yielded both superior accuracy ($r=0.64$) and interpretable attention maps. This finding has direct implications for applying CDT-II to new systems: researchers should prioritize genes whose perturbation effects are reproducible across independent experiments. The elimination of RNA and protein language model dependencies—a key design change from CDT v18—enables this flexibility by allowing researchers to work directly with raw per-cell expression rather than fixed embeddings.
@@
A notable feature of CDT-II is its architectural simplicity. The model uses only standard transformer components—self-attention, cross-attention, and feed-forward layers—without specialized modules, custom loss functions, or complex training procedures. This simplicity is deliberate: by constraining the architecture to mirror the central dogma rather than adding complexity, each attention mechanism retains a clear biological interpretation. In contrast, more complex architectures often sacrifice interpretability for marginal performance gains. CDT-II demonstrates that biological insight can emerge from standard components when they are structured to reflect biological organization.
@@
Several limitations should be noted. The overall per-cell validation correlation ($r=0.64$), while below unity, should be evaluated against the empirical reproducibility ceiling of CRISPRi experiments. The cross-study correlation between Morris et al. STING-seq and Gasperini et al. CRISPRi screens is approximately $r\approx 0.79$ at the pseudo-bulk level8,10,12; CDT-II’s per-cell $r=0.64$ thus exceeds 80% of this experimental ceiling, and its pseudo-bulk correlation (mean $r=0.84$) surpasses the inter-experimental reproducibility. Furthermore, Morris et al. reported that only $\sim$56% of GFI1B trans-targets replicated between these two independent studies10,12, confirming that the remaining prediction gap largely reflects irreducible experimental variability rather than model limitations. While in silico knockdown experiments—setting gene expression to zero—yielded weak correlations ($r\approx 0.07$), this reflects the challenge of extrapolating far outside the training distribution rather than a failure to learn regulatory relationships. In contrast, gradient-based attribution ($\partial\text{output}_{j}/\partial\text{input}_{i}$), which measures local sensitivity within the training distribution, achieved $r=0.83$ correlation with experimental effect sizes. This demonstrates that CDT-II accurately captures regulatory relationships when queried appropriately, and suggests that gradient-based virtual screening—predicting the direction and relative magnitude of perturbation effects—may be more practical than simulating complete knockdowns. The 2,361-gene set excludes some known GFI1B targets (INSIG1, NPM1, BNIP3), and all experiments were conducted in K562 cells from a single dataset. Finally, attention weights require careful interpretation20,21; we mitigate this concern through external validation (6.6-fold enrichment) rather than relying on attention as post-hoc explanation, but attention should not be interpreted as causal without experimental follow-up.
@@
Looking ahead, CDT-II opens several avenues for biological discovery. First, gradient-based virtual screening ($r=0.83$ with experimental effects) could enable computational prioritization of perturbation experiments: by computing input gradients for any genomic locus in a trained model, researchers could rank regulatory targets by predicted effect magnitude before committing experimental resources. Second, extension to multiple cell types would test whether regulatory networks are cell-type-specific or conserved, potentially revealing shared regulatory modules across lineages. Third, integration of additional modalities—scATAC-seq for chromatin accessibility, DNA methylation for epigenetic state, or spatial transcriptomics for tissue context—could further sharpen the microscope’s resolution by providing the model with richer cellular context. Fourth, CDT-II’s modular architecture allows the DNA embedding component to be upgraded as genomic foundation models improve; Enformer could be replaced with more recent models such as AlphaGenome22 or Evo23, potentially improving the resolution of DNA-mediated regulatory predictions without retraining the entire system. Finally, because CDT-II learns from perturbation data rather than observational correlations, it may capture causal regulatory relationships that could inform therapeutic target identification in disease contexts.
@@
Primary CRISPRi dataset. We used the STING-seq v2 dataset10 (GEO accession via original publication), which profiled 60,505 K562 cells using single-cell RNA sequencing combined with CRISPR interference9. Perturbations targeted 447 genomic loci: 27 transcription start sites (TSS) and 420 single-nucleotide polymorphism (SNP) loci. Each cell received a single guide RNA targeting one locus. Non-targeting control (NTC) cells (n=45,226n=45{,}226) served as the unperturbed baseline.
Gene set curation. The 2,361-gene set was derived by intersecting genes detected as expressed in both an independent K562 CRISPRi screen12 and the primary STING-seq dataset10, further requiring presence in the scGPT vocabulary and ProteomeLM database to maintain compatibility with CDT v18 for direct architectural comparison. This cross-dataset reproducibility filter yielded 2,360 genes. GFI1B was added as a 2,361st gene based on its established role as a master transcriptional regulator with hundreds of characterized trans-regulatory targets across the transcriptome in the primary dataset10, providing a critical test case for unsupervised network recovery.
@@
where $x_{cg}$ is the CPM-normalized expression of gene $g$ in cell $c$, and $\bar{x}_{g}^{\text{NTC}}$ is the mean CPM expression of gene $g$ across all 45,226 non-targeting control cells. This per-cell computation preserves individual cell-state variation rather than collapsing to pseudo-bulk effect sizes.
DNA embeddings. Enformer3 pre-computed embeddings of shape [896, 3,072] were generated for each perturbation locus, centered on either the TSS or SNP position, representing a 196-kilobase genomic window at 128-base-pair resolution.
@@
The CDT-II architecture (Fig. 1) follows the design principles established in CDT v18, with key modifications for raw expression input. The RawExpressionEncoder combines learned gene identity embeddings with projected log1p(CPM) expression values to produce gene-level representations. Enformer DNA embeddings are projected from 3,072 to 512 dimensions. All attention operations use standard scaled dot-product attention11 with 8 heads, embedding dimension $d=512$, FFN hidden dimension 4d=2,0484d=2{,}048, GELU activation, and dropout $p=0.3$. DNA representations pass through 2 self-attention layers; RNA representations pass through 1 self-attention layer followed by 1 cross-attention layer (RNA queries, DNA keys/values). The Virtual Cell Embedder8 uses 4-head attention pooling to compress each modality into a fixed-dimensional vector, which are concatenated and fused via MLP. A two-layer task head predicts log2 fold changes for all 2,361 genes. Total parameters: $\sim$21 million.
@@
where $\hat{y}_{cg}$ is the predicted log2 fold change for gene $g$ in cell $c$, $y_{cg}$ is the observed value, G=2,361G=2{,}361, and $\mathcal{B}$ is a mini-batch. We used the AdamW18 optimizer (learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-5}$) with ReduceLROnPlateau scheduling (factor 0.5, patience 10 epochs, monitoring validation loss). Gradient norms were clipped to a maximum of 1.0. Batch size was 64. Training was performed on a single NVIDIA A100 GPU (40 GB).
@@
Community detection. Louvain community detection14 was applied to identify modules of co-regulated genes. For RNA self-attention, this yielded 5 communities; for cross-attention similarity, 53 communities (reflecting finer-grained structure).
GO enrichment. Gene Ontology15 and pathway enrichment was performed using the Enrichr19 API against Gene Ontology Molecular Function, KEGG17, and Reactome16 databases. $P$ values were corrected for multiple testing.
@@
Two principal training configurations were compared. The full gene set configuration used 9,335 genes detected in at least 10% of cells in the primary dataset10 (excluding mitochondrial, ribosomal, and hemoglobin genes), applying a single-dataset expression threshold. The filtered gene set configuration used 2,361 genes derived from a cross-dataset reproducibility filter (intersection of two independent CRISPRi screens10,12) plus GFI1B. Both configurations used identical cell splits (13,620 training, 2,037 validation) and architecture. Learning curves were compared on the same validation cell set using Pearson $r$.
@@
CDT-II source code is available at https://github.com/nobusama/CDT2. Enformer embeddings, processed training data, and trained model weights are available at https://huggingface.co/datasets/nobusama17/CDT2-data. The STING-seq v2 dataset is available from the original publication10. CDT v1 is described in ref. 8.
@@ Acknowledgements
The author thanks the developers of Enformer, the Morris laboratory for making the STING-seq dataset publicly available, and the open-source communities behind PyTorch and related tools.

### v2

@@ Discussion
We note that Enformer’s pre-training on epigenomic tracks across many cell types means the initial DNA embeddings encode general regulatory element positions based on sequence alone—Enformer receives no cell-type information and produces identical embeddings regardless of cellular context. These cell-type-agnostic embeddings are then substantially transformed before cross-attention—through a learned projection layer and two DNA self-attention layers, all trained on K562 perturbation prediction—so the cross-attention maps reflect CDT-II’s cell-type-specific, task-specific understanding rather than Enformer’s general-purpose representations. The key finding is that this multi-layer transformation selects the gene-specific regulatory elements relevant to each perturbation target in K562, transforming Enformer’s cell-type-agnostic genomic representation into an interpretable, gene-resolved regulatory map that aligns with K562 ENCODE annotations20. More broadly, CDT-II is designed as an integration platform: its goal is not to outperform any individual component, but to combine pre-trained representations with experimental perturbation data and extract interpretable regulatory structure that neither source can provide alone. The DNA embedding module is deliberately modular—Enformer can be replaced with newer models such as AlphaGenome21 or Evo22—so the relevant question is not how much information originates from any single component, but whether the integrated system produces biologically meaningful outputs.
The prominence of CTCF in cross-attention maps has implications for understanding three-dimensional chromatin organization. CTCF is a primary architectural protein that establishes topologically associating domain (TAD) boundaries and chromatin loop anchors, compartmentalizing the genome into regulatory neighborhoods23. The finding that CTCF-associated bins receive the highest attention weights across all five held-out genes (Cohen’s $d=2.4$) and show the most consistent gene-specificity signal suggests that CDT-II may be learning aspects of three-dimensional regulatory structure from one-dimensional sequence information alone. This is particularly striking because the model receives no explicit chromatin conformation data—neither Hi-C contact maps nor TAD annotations. Future integration of chromatin conformation data could further sharpen this aspect of the AI microscope’s resolution, potentially enabling direct visualization of gene-specific regulatory neighborhoods.
The ablation study establishes a practically important principle: gene set quality, determined by cross-dataset reproducibility filtering, governs the resolution of the AI microscope. Training with 9,335 genes produced inferior predictions ($r=0.37$), while 2,361 cross-dataset validated genes yielded both superior accuracy ($r=0.64$) and interpretable attention maps. This finding has direct implications for applying CDT-II to new systems: researchers should prioritize genes whose perturbation effects are reproducible across independent experiments. The elimination of RNA and protein language model dependencies—a key design change from CDT v110—enables this flexibility by allowing researchers to work directly with raw per-cell expression rather than fixed embeddings.
@@
A notable feature of CDT-II is its architectural simplicity. The model uses only standard transformer components—self-attention, cross-attention, and feed-forward layers—without specialized modules, custom loss functions, or complex training procedures. This simplicity is deliberate: by constraining the architecture to mirror the central dogma rather than adding complexity, each attention mechanism retains a clear biological interpretation. In contrast, more complex architectures often sacrifice interpretability for marginal performance gains. CDT-II demonstrates that biological insight can emerge from standard components when they are structured to reflect biological organization. This approach parallels the interpretability-by-design principle demonstrated by BPNet24 for transcription factor binding, extended here to multi-modal regulatory network inference.
@@
Several limitations should be noted. CDT-II’s pseudo-bulk correlation (mean $r=0.84$) surpasses the inter-experimental reproducibility ceiling ($r\approx 0.79$ between Morris et al. and Gasperini et al. screens10,12,14), suggesting that the remaining per-cell prediction gap ($r=0.64$) largely reflects irreducible single-cell stochasticity and inter-experimental variability rather than model limitations. While in silico knockdown experiments—setting gene expression to zero—yielded weak correlations ($r\approx 0.07$), this reflects the challenge of extrapolating far outside the training distribution rather than a failure to learn regulatory relationships. In contrast, gradient-based attribution ($\partial\text{output}_{j}/\partial\text{input}_{i}$), which measures local sensitivity within the training distribution, achieved $r=0.83$ correlation with experimental effect sizes. This demonstrates that CDT-II accurately captures regulatory relationships when queried appropriately, and suggests that gradient-based virtual screening—predicting the direction and relative magnitude of perturbation effects—may be more practical than simulating complete knockdowns. The 2,361-gene set excludes some known GFI1B targets (INSIG1, NPM1, BNIP3), and all experiments were conducted in K562 cells from a single dataset. Finally, attention weights require careful interpretation25,26; we mitigate this concern through external validation (6.6-fold enrichment) rather than relying on attention as post-hoc explanation, but attention should not be interpreted as causal without experimental follow-up. We note that direct comparison with perturbation prediction methods such as GEARS6,27 and CPA7 is not applicable: those methods aim to predict the effects of unseen perturbations without experimental data, whereas CDT-II uses experimental perturbation data to reveal interpretable regulatory structure—fundamentally different objectives requiring different evaluation criteria. Similarly, network inference methods such as SCENIC28 and CellOracle29 reconstruct regulatory networks from observational data; CDT-II complements these by deriving regulatory structure from perturbation responses, grounding network edges in causal experimental data.
A distinctive feature of the AI microscope framework is that its outputs generate experimentally testable hypotheses. The CTCF binding sites and enhancer elements identified by cross-attention (Fig. 6) represent specific genomic loci that can be validated through targeted CRISPRi experiments, and the gene regulatory edges recovered by RNA self-attention (Fig. 5) predict specific regulatory relationships testable by individual gene knockdowns. This creates a feedback cycle: experimental validation of attention-derived hypotheses not only tests the model but also generates new training data that can improve its resolution. As genomic foundation models continue to advance, CDT-II provides a means for experimentalists to interrogate, validate, and refine these models’ representations through their own data—closing the loop between computational prediction and experimental biology.
@@
Looking ahead, CDT-II opens several avenues for biological discovery. First, gradient-based virtual screening ($r=0.83$ with experimental effects) could enable computational prioritization of perturbation experiments: by computing input gradients for any genomic locus in a trained model, researchers could rank regulatory targets by predicted effect magnitude before committing experimental resources. Second, the growing availability of genome-scale perturbation atlases30 across multiple cell types provides an immediate path for testing whether regulatory networks are cell-type-specific or conserved, potentially revealing shared regulatory modules across lineages. Third, integration of additional modalities—scATAC-seq for chromatin accessibility, DNA methylation for epigenetic state, or spatial transcriptomics for tissue context—could further sharpen the microscope’s resolution by providing the model with richer cellular context. Fourth, CDT-II’s modular architecture allows the DNA embedding component to be upgraded as genomic foundation models improve; Enformer could be replaced with more recent models such as AlphaGenome21 or Evo22, potentially improving the resolution of DNA-mediated regulatory predictions without retraining the entire system. Finally, because CDT-II learns from perturbation data rather than observational correlations, it may capture causal regulatory relationships that could inform therapeutic target identification in disease contexts.
@@
Primary CRISPRi dataset. We used the STING-seq v2 dataset12 (GEO accession GSE171452), which profiled 60,505 K562 cells using single-cell RNA sequencing combined with CRISPR interference11,31. Perturbations targeted 447 genomic loci: 27 transcription start sites (TSS) and 420 single-nucleotide polymorphism (SNP) loci. Because the screen is highly multiplexed (mean 93 guides per cell), most cells lacked sufficient guide RNA signal for unambiguous assignment; cells were retained only if exactly one target gene had UMI $\geq 50$ with no competing signal, yielding 10,328 assignable cells (8,250 TSS-perturbed and 2,078 NTC) from the original 60,505. Non-targeting control (NTC) cells (n=2,078n=2{,}078) served as the unperturbed baseline. An additional 7,407 SNP-perturbed cells (420 loci) were used with guide assignments from the original publication12.
Gene set curation. The 2,361-gene set was derived by intersecting genes detected as expressed in both an independent K562 CRISPRi screen14 and the primary STING-seq dataset12. This cross-dataset reproducibility filter yielded 2,360 genes. GFI1B was added as a 2,361st gene based on its established role as a master transcriptional regulator with hundreds of characterized trans-regulatory targets across the transcriptome in the primary dataset12, providing a critical test case for unsupervised network recovery.
@@
where $x_{cg}$ is the CPM-normalized expression of gene $g$ in cell $c$, and $\bar{x}_{g}^{\text{NTC}}$ is the mean CPM expression of gene $g$ across all 2,078 non-targeting control cells. This per-cell computation preserves individual cell-state variation rather than collapsing to pseudo-bulk effect sizes.
DNA embeddings. Pre-computed Enformer3 embeddings of shape [896, 3,072] were generated for each perturbation locus, centered on either the TSS or SNP position, spanning approximately 115 kilobases at 128-base-pair resolution.
@@
The CDT-II architecture (Fig. 1) follows the design principles established in CDT v110, with key modifications for raw expression input. The RawExpressionEncoder combines learned gene identity embeddings with projected log1p(CPM) expression values to produce gene-level representations. Enformer DNA embeddings are projected from 3,072 to 512 dimensions. All attention operations use standard scaled dot-product attention13 with 8 heads, embedding dimension $d=512$, FFN hidden dimension 4d=2,0484d=2{,}048, GELU activation, and dropout $p=0.3$. DNA representations pass through 2 self-attention layers; RNA representations pass through 1 self-attention layer followed by 1 cross-attention layer (RNA queries, DNA keys/values). The Virtual Cell Embedder10 uses 4-head attention pooling to compress each modality into a fixed-dimensional vector, which are concatenated and fused via MLP. A two-layer task head predicts log2 fold changes for all 2,361 genes. Total parameters: $\sim$21 million.
@@
where $\hat{y}_{cg}$ is the predicted log2 fold change for gene $g$ in cell $c$, $y_{cg}$ is the observed value, G=2,361G=2{,}361, and $\mathcal{B}$ is a mini-batch. We used the AdamW32 optimizer (learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-5}$) with ReduceLROnPlateau scheduling (factor 0.5, patience 10 epochs, monitoring validation loss). Gradient norms were clipped to a maximum of 1.0. Batch size was 64. Training was performed on a single NVIDIA A100 GPU (40 GB).
@@
ENCODE enrichment analysis
Cross-attention profiles were compared against ENCODE K562 regulatory annotations20 to assess whether attention autonomously focuses on known regulatory elements. Five ENCODE K562 peak-call datasets were used: CTCF ChIP-seq (ENCFF769AUF), H3K27ac ChIP-seq (ENCFF864OSZ), H3K4me1 ChIP-seq (ENCFF135ZLM), H3K4me3 ChIP-seq (ENCFF313FYW), and DNase-seq (ENCFF422QRZ).
Enrichment testing. For each held-out gene, the mean cross-attention profile (averaged across validation cells and 8 attention heads) was computed over the 896 genomic bins. Bins in the top 10% of attention weight were designated as “high-attention.” Each ENCODE peak file was intersected with the 896-bin coordinate system, and Fisher’s exact tests were computed for each gene–mark combination (25 tests: 5 genes $\times$ 5 marks). Odds ratios were Haldane-corrected (adding 0.5 to each cell of the $2\times 2$ table) to handle zero cells and provide finite 95% confidence intervals. All 23 significant $P$ values were below $1\times 10^{-3}$, well within the Bonferroni-corrected threshold of $0.05/25=0.002$.
Circular permutation test. To control for spatial autocorrelation in both attention profiles and ENCODE peaks, we performed circular permutation tests (n=1,000n=1{,}000). In each permutation, the attention profile was circularly shifted by a random offset along the genomic axis, preserving the autocorrelation structure of the attention signal while disrupting its alignment with ENCODE peaks. The observed enrichment was compared against the null distribution to obtain permutation $P$ values.
Bin classification and effect sizes. Each genomic bin was classified into one of four categories based on ENCODE annotations: promoter (H3K4me3+), active enhancer (H3K27ac+ & H3K4me3-), CTCF-only (CTCF+ & H3K4me3- & H3K27ac-), or unannotated (no peaks). Attention weights were compared across bin classes using Kruskal–Wallis tests. Pairwise effect sizes between each annotated class and unannotated bins were quantified using Cohen’s $d$.
@@
Community detection. Louvain community detection16 (resolution parameter = 1.0, the default) was applied to identify modules of co-regulated genes. For RNA self-attention, this yielded 5 communities; for cross-attention similarity, 53 communities (reflecting finer-grained structure).
GO enrichment. Gene Ontology17 and pathway enrichment was performed using the Enrichr33 API against Gene Ontology Molecular Function, KEGG19, and Reactome18 databases. $P$ values were corrected for multiple testing using the Benjamini–Hochberg procedure.
@@
Two principal training configurations were compared. The full gene set configuration used 9,335 genes detected in at least 10% of cells in the primary dataset12 (excluding mitochondrial, ribosomal, and hemoglobin genes), applying a single-dataset expression threshold. The filtered gene set configuration used 2,361 genes derived from a cross-dataset reproducibility filter (intersection of two independent CRISPRi screens12,14) plus GFI1B. Both configurations used identical cell splits (13,620 training, 2,037 validation), architecture template, and hyperparameters; parameter counts differed due to gene set size ($\sim$21 million vs $\sim$54 million). Learning curves were compared on the same validation cell set using Pearson $r$.
@@
CDT-II source code is available at https://github.com/nobusama/CDT2. Enformer embeddings, processed training data, and trained model weights are available at https://huggingface.co/datasets/nobusama17/CDT2-data. The STING-seq v2 dataset is available from the original publication12. CDT v1 is described in ref. 10.
@@ Acknowledgements
The author thanks the developers of Enformer, the Morris laboratory for making the STING-seq dataset publicly available, and the open-source communities behind PyTorch and related tools. Claude (Anthropic) was used for manuscript proofreading and editing. The author assumes full responsibility for the accuracy and integrity of all content in this manuscript.
@@
Supplementary Information
Figure 7: Supplementary Figure 1: Statistical robustness of ENCODE enrichment. (a) Circular permutation test (n=1,000n=1{,}000): observed enrichments (red lines) versus null distributions for GFI1B across three marks (DNase, CTCF, H3K4me3), confirming robustness against spatial autocorrelation artifacts. (b) Threshold robustness: enrichment odds ratios remain significant across attention thresholds from top 5% to top 20%, indicating that results are not dependent on a single arbitrary cutoff.